Clear Vision Entertainment, Inc. is a digital satellite television provider that charges $35 per month for service. The company plans to increase marketing expenditures aimed at customer retention next year, but wants to determine CLV before making this decision. Variable Costs before marketing are $5 per account per month. Per year, the company spends $12 per account in retention costs spread equally across the year. Clear Vision's annual attrition rate is 20% with a monthly discount rate of 1%.

  1. What is the monthly contribution margin ($) per customer account (not including retention costs)?

  2. What are the monthly marketing retention costs per customer account?

  3. What is the monthly retention rate?
Answer #1

Monthly contribution margin ($) = Sevice charge - Variable cost

= 35 - 5

= $ 30

Monthly marketing retention cost per customer account = Annual retention cost per account / 12

= $ 12 / 12

= $ 1

Monthly retention rate = (1 - Annual attrition rate)

= (1-20%)

= 80 %

CLV = Margin * Retention rate / (1 + discount rate - retention rate)

= (30-1)*80%/(1+1%-80%)

= $ 110.5